Help!! (managerial accounting)

The Sweetwater Candy Company would like to buy a new machine that would automatically “dip” chocolates. The dipping operation currently is done largely by hand. The machine the company is considering costs $180,000. The manufacturer estimates that the machine would be usable for five years but would require the replacement of several key parts at the end of the third year. These parts would cost $9,900, including installation. After five years, the machine could be sold for $5,000.

 

 The company estimates that the cost to operate the machine will be $7,900 per year. The present method of dipping chocolates costs $39,000 per year. In addition to reducing costs, the new machine will increase production by 6,000 boxes of chocolates per year. The company realizes a contribution margin of $1.30 per box. A 15% rate of return is required on all investments.

 

Click here to view Exhibit 12B-1 and Exhibit 12B-2, to determine the appropriate discount factor(s) using tables.

 

Required:

1. What are the annual net cash inflows that will be provided by the new dipping machine?

2. Compute the new machine’s net present value.
Answer #1

a.) What are the annual net cash inflows that will be provided by the new dipping

     machine?

     Annual Savings in Operating Cost: $311,000= 39,000- 7,900

     (Add) Contribution Margin: $7,800= 6,000(1.30)

    Net Annual Cash Inflows: $38,900


b.) Compute the new machine’s net present value.

     Cash Outflows: Cost of Machine: ($180,000)

                              Cost of Replacement Parts: $6,514.20= 9,900(0.658)

     Present Value of Cash Outflows: $186,514.20


     Cash Inflows: Annual Cash Inflow: $130,392.80= 38,900(3.352)

                            Salvage Value: $2,485= 5,000(0.497)

     Present Value of Cash Inflows: $132,877.80


     New Present Value: ($53,636.40)

source: Connect Assignment